1. The town of Springfield has two steel producers that pollute the air with Sulfur Dioxide (SO). Their supply curves are sho i need question b please
Answer #1

Initially we have, Supply of firem A : P = QA supply of firm B : P = QB Market Supply : P = pm 2 [Qm= QATQ The p in represeSo based on i representing social o Ps = QA + 2 I opg = QB + 2 our supply functions cost will be, [Firma] [firm o] + 2 om A P- Put Om = 18 in Market Supply equation to get socially optimal price , Ps = Qm +2 Ps = 18 + 2 - Ps = 6+2=8 socially optimal